4. (a) A television repairman finds that the time spent on his jobs has an exponential

distribution with a mean of 30 minutes. If he repairs sets in the order in which they

come in, and if arrival of sets follows a Poission distribution approximately with an

average rate of 10 per 8 hours day, what is the repairman’s expected idle time each

day, How many jobs are ahead of the average set just brought in?


1
Expert's answer
2021-04-28T08:49:50-0400

Here"\lambda=\frac{10}{8×60}=\frac{1}{48}set per minute"


But "\mu=\frac{1}{30} set/minute"

Prob.that their is no unit in the system PO=1-"\frac{\lambda}{\mu}"


="1-\frac{5}{8}=\frac{3}{8}"

Repairmans's expected idle time in 8 hours a day:

=nPO="8×\frac{3}{8}" =3hours


Expected average n.o. of jobs or average n.o.of Tv sets in the system

LS="\frac{\lambda}{\mu-\lambda}"

But "\mu-\lambda"

="\frac{1}{30}-\frac{1}{48}=\frac{1}{80}"

Substuting into the formula


"=\frac {1}{48}×\frac{80}{1}=\frac{5}{3}jobs"